OpenStudy (anonymous):

Which of the following are true statements.   The lengths 14, 24 and 26 can not be sides of a right triangle.  The lengths 30, 72, and 78 can not be sides of a right triangle.   The lengths 14, 24 and 26 can be sides of a right triangle.  The lengths 30, 72, and 78 can be sides of a right triangle.  .The lengths 14, 24 and 26 can not be sides of a right triangle.  The lengths 30, 72, and 78 can be sides of a right triangle.  The lengths 14, 24 and 26 can be sides of a right triangle.  The lengths 30, 72, and 78 can not be sides of a right triangle.

OpenStudy (anonymous):

This isn't too difficult, it just requires a lot of repetition. You need to test each set of sides. Pythagoreans theorem should prove useful. a^2 + b^2 = c^2 a and b are your shortest sides and c is the longest side. If it's true then it's a right triangle, if it's false then it's not a right triangle. For example, your first question. 14^2 +24^2 =26^2 This translates to 196+576=676 Is that true? 772 is not equal to 676 so your first line is NOT a right triangle. Is that enough information? :)